OpenStudy (anonymous):

Which of the following is equivalent to the proportion 2 over 3 equals 3 over y? A. 3y = (2)(3) B. 2y = (2)(3) C. 2y = (3)(3) D. 3y = (3)(3)

OpenStudy (alienium):

2/3 = 3/x x = 9/2 :)

OpenStudy (alienium):

yes